# Prompting Large Language Models for Zero-shot Essay Scoring via Multi-trait Specialization

### v1

@@ Abstract
Advances in automated essay scoring (AES) have traditionally relied on labeled essays, requiring tremendous cost and expertise for their acquisition. Recently, large language models (LLMs) have achieved great success in various tasks, but their potential is less explored in AES. In this paper, we propose Multi Trait Specialization (MTS), a zero-shot prompting framework to elicit essay scoring capabilities in LLMs. Specifically, we leverage ChatGPT to decompose writing proficiency into distinct traits and generate scoring criteria for each trait. Then, an LLM is prompted to extract trait scores from several conversational rounds, each round scoring one of the traits based on the scoring criteria. Finally, we derive the overall score via trait averaging and min-max scaling. Experimental results on two benchmark datasets demonstrate that MTS consistently outperforms straightforward prompting (Vanilla) in average QWK across all LLMs and datasets, with maximum gains of 0.437 on TOEFL11 and 0.355 on ASAP. Additionally, with the help of MTS, the small-sized Llama2-13b-chat substantially outperforms ChatGPT, facilitating an effective deployment in real applications.
Prompting Large Language Models for Zero-shot Essay Scoring
via Multi-trait Specialization
[…]
@@ 1 Introduction
Automated essay scoring (AES) aims at evaluating and scoring essays with machine learning and natural language processing (Dikli, 2006). AES is a promising alternative to costly and laborious human assessment, greatly resolving rater fatigue and inter-rater inconsistency. AES systems have been widely deployed in classroom settings (Dikli and Bleyle, 2014) and high-stakes tests such as TOEFL (Attali and Burstein, 2006).
@@
Figure 1: Comparison of our MTS zero-shot prompting framework for AES and Vanilla baseline across different types of LLMs and datasets, based on average QWK.
@@
Previous studies highly matched human ratings via developing supervised models tailored to specific prompts (Yannakoudakis et al., 2011; Taghipour and Ng, 2016; Dong et al., 2017), assuming essays from the train and test sets belong to the same prompt. However, prompt-specific models struggled when confronted with essays written for unseen prompts (Jin et al., 2018; Cozma et al., 2018). Hence AES is progressing towards reflecting more real-world scenarios, exemplified by cross-prompt approaches (Cao et al., 2020; Jiang et al., 2023b) which reinforce domain transferability of the supervised models.
@@
Recently, advances in large language models (LLMs) (Brown et al., 2020; Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023) have led to a paradigm shift in which LLMs excel across a wide range of downstream tasks via zero-shot or few-shot instructions (Yuan et al., 2023; Zhang et al., 2023a). In many cases, careful prompt design plays a crucial role in unlocking LLMs’ potential. For instance, chain-of-thought (CoT) prompting (Wei et al., 2022; Kojima et al., 2022) dramatically improves LLMs’ performance on complex reasoning benchmarks by externalizing the reasoning process.
@@
The development of LLM-based chatbot aligned with human preferences (Ouyang et al., 2022; Rafailov et al., 2023) has given rise to zero-shot AES, allowing us to move beyond the cross-prompt setting. However, leveraging LLMs for zero-shot AES is less explored, in contrast to proliferating studies harnessing LLMs to serve as an evaluation metric for machine-generated text (Chiang and Lee, 2023; Zheng et al., 2023; Liu et al., 2023). Initial works of zero-shot AES (Mizumoto and Eguchi, 2023; Yancey et al., 2023) prompt LLMs to assign the overall score within a single step, which demonstrate suboptimal agreement with human raters.
@@
In this paper, we present MTS (Multi Trait Specialization), a zero-shot prompting framework to elicit essay scoring capabilities in LLMs, partially inspired by the supervised models with trait-attention mechanism (Ridley et al., 2021; Do et al., 2023). Specifically, we exploit ChatGPT (OpenAI, 2022) to decompose the writing quality into multiple traits and generate scoring criteria for each trait. Next, an LLM engages in several rounds of conversation, each round evaluating with respect to one of the traits. During the conversation, the LLM is instructed to retrieve quotes to provide faithful evaluation on the essay, then assign a score based on the given scoring criteria. Finally, the overall score is derived by averaging and min-max scaling the trait scores, in combination with the outlier clipping mechanism.
@@
We evaluate MTS on ASAP (Hamner et al., 2012) and TOEFL11 (Blanchard et al., 2013) with different LLMs, including ChatGPT, Llama 2 (Touvron et al., 2023) and Mistral 7b (Jiang et al., 2023a). We take the Vanilla approach as a baseline, which prompts LLM to directly assign an overall score along with CoT. As illustrated in Figure 1, MTS consistently outperforms Vanilla in average Quadratic Weighted Kappa (QWK) across all combinations of LLMs and datasets, with maximum gains of 0.437 ($0.025\rightarrow 0.462$) on TOEFL11 and 0.355 ($0.205\rightarrow 0.560$) on ASAP. In addition, the small-sized Llama2-13b-chat substantially surpasses ChatGPT with the help of MTS, enabling a more effective deployment. Moreover, MTS trails supervised SOTA in average QWK only by 0.171 on TOEFL11 and 0.257 on ASAP, which offers a strong zero-shot alternative to supervised models.
@@
  • We leverage ChatGPT to decompose the AES task into simpler subtasks of assigning a score with respect to a specific trait.
  • We devise Quote Retrieval and Scoring mechanism which facilitates generating faithful evaluations to boost the accuracy of scoring.
  • We utilize min-max scaling with outlier clipping which effectively maps predicted scores to arbitrary target score ranges.
  • Our proposed method achieves promising results, significantly exceeding Vanilla, while also outperforming ChatGPT with the small-sized Llama2-13b-chat.
Figure 2: Instruction for multi trait decomposition used in ASAP. The contents to be filled are highlighted in red. See Appendix A for the instructions used in TOEFL11.
@@ 2 Method
We start by formalizing the definition of zero-shot AES. Given a dataset consisting of unlabelled essays $\mathcal{D}=\{x^{(i)}\}_{i=1}^{N_{\mathcal{D}}}$, the goal is to output prediction $\hat{y}^{(i)}$ for every essay $x^{(i)}$ from a set of predefined scores $\mathcal{Y}$ where $\hat{y}^{(i)}\in\mathcal{Y}$. We achieve this goal via zero-shot prompting without parameter tuning.
@@ 2.1 Multi Trait Decomposition
We address this issue by decomposing the writing proficiency into several key traits and defining trait-specific scoring criteria which will be fixed during the assessment (details in Section 2.2), contributing to a consistent scoring behavior of LLM. We automate these preliminaries by leveraging the power of LLM. As shown in Figure 2, we instruct ChatGPT to condense the rubric guidelines used by human raters into several key traits of writing quality and generate scoring criteria for each trait. This procedure ensures the trait-specific scoring criteria could be isolated from the rubric guidelines in which multiple aspects of writing are mentioned in a mixed way. An example of the generation result is depicted in the left part of Figure 3 (see more results in Appendix B). Note that the generation result varies across different rubric guidelines.
@@ 2.2 Trait Specialization
In Step 1, leveraging trait-specific conversations not only simplifies the AES task as a form of subproblem decomposition but also prevents evaluations on each trait from being influenced by the other traits.
In Step 2 and 3, we present a new mechanism where a quote retrieval task is appended before the main scoring task, as illustrated in the middle part of Figure 3. The quote retrieval task allows LLM to adhere to the details of the essay and avoid producing generic evaluation, whereas the scoring task puts emphasis on the mapping between the evaluation on the quotes and the scoring criteria.
@@ 2.3 Trait Aggregation and Scaling
Parsing the output of Section 2.2 yields trait scores $\{\hat{y}_{j}^{(i)}\}_{j=1}^{N_{T}}$ for an essay $x^{(i)}$, where $N_{T}$ represents the number of predefined traits. The trait scores should be transformed to a final prediction $\hat{y}^{(i)}$ that falls under the target score range which may vary greatly across different prompts (as in Table 1).
Based on the independent trait scores, we propose a new trait aggregation and scaling method. First, the trait scores are aggregated by taking their average, and the outliers among the averaged scores are clipped using Q1 and Q3, i.e. the first and the third quartiles of the averaged scores, that is, the clipped score $\hat{y}_{agg}^{(i)}$ is computed as follows:
@@
where $\hat{y}_{min}=\min\limits_{i}\hat{y}_{agg}^{(i)}$ and $\hat{y}_{max}=\max\limits_{i}\hat{y}_{agg}^{(i)}$.
In this way, the clipping alleviates the sensitivity of min-max scaling to the outliers.
@@ 3.1 LLMs
To verify the effectiveness of our proposed method is universal across LLMs, we choose different types of LLMs which are not variants of one another: ChatGPT, Llama 2 and Mistral 7b. In detail, we use their instruction-tuned models which are gpt-3.5-turbo-0613, Llama2-7b-chat, Llama2-13b-chat and Mistral-7B-Instruct-v0.2. The temperature is set to $0.1$ for all LLMs and the repetition penalty is set to $1.1$ for all LLMs but ChatGPT. Other hyperparameters for sampling follow the defaults. Experiments are run once with a fixed random seed considering limitations of computational resources.
@@ 3.2 Datasets and Evaluation Metric
We conduct experiments on two datasets, ASAP [^1] https://www.kaggle.com/c/asap-aes/data and TOEFL11. ASAP (Automated Student Assessment Prize) comprises 12,978 essays provided by students spanning in grade level from 7 to 10. The essays were written in response to 8 prompts of varying genres and score ranges. TOEFL11 consists of 12,100 essays written by test takers of TOEFL iBT (2006-2007) which measures the academic English proficiency (low/medium/high) of non-native English speakers. The statistics of the two datasets are shown in Table 1.
For ASAP, 10% of the data from each prompt are sampled for testing. There is no significant difference in the average essay score between the sample and the population in all prompts (Z-test, $\alpha=0.05$). To our knowledge, we are the first to benchmark zero-shot performance of LLMs on the widely used ASAP dataset. For TOEFL11 dataset, we adopt the same test split from the 2013 Native Language Identification Shared Task (Tetreault et al., 2013), which consists of 1,100 essays collected from 8 prompts.
We use Quadratic Weighted Kappa (QWK) to measure the agreement between groundtruth scores and predicted scores, which is widely used in previous AES research (Taghipour and Ng, 2016; Dong et al., 2017; Cao et al., 2020; Xie et al., 2022).
@@
Dataset Category LLM Method P1 P2 P3 P4 P5 P6 P7 P8 Avg.
ASAP fine-tuned N/A NPCR 0.856 0.750 0.756 0.851 0.847 0.858 0.838 0.779 0.817
zero-shot ChatGPT Vanilla 0.032 0.220 0.476 0.597 0.479 0.637 0.289 0.527 0.407
@@
TOEFL11 fine-tuned N/A NPCR 0.750 0.706 0.806 0.725 0.802 0.748 0.731 0.798 0.758
zero-shot ChatGPT Vanilla 0.215 0.240 0.337 0.332 0.227 0.306 0.237 0.306 0.275
@@
Llama2-7b-chat Vanilla $0$0.007 0.026 -0.006 0.041 0.015 0.111 0.020 0.025
@@
Llama2-13b-chat Vanilla 0.196 0.156 0.285 0.268 0.165 0.329 0.249 0.257 0.238
@@
Mistral-7b-instruct Vanilla 0.486 0.259 0.355 0.344 0.431 0.456 0.286 0.383 0.375
@@
Table 2: Evaluation results in QWK on ASAP and TOEFL11. †: The data are from the original paper; ‡: Our implementation. Negative value indicates that the predictions are worse than random.
@@ 3.3 Implementation Details
Rubric Guidelines. As outlined in Section 2.1, ChatGPT generates scoring criteria based on the rubric guidelines for human raters. In ASAP, each prompt has rubric guidelines distinct from the other prompts, resulting in the scoring criteria varying across the prompts. In contrast, TOEFL adopts unified rubric guidelines, leading to the scoring criteria identical across all prompts. Following Mizumoto and Eguchi (2023), when generating the scoring criteria, we replace TOEFL Independent Writing Rubrics with IELTS Task2 Writing Band Descriptor since the IELTS rubric (10-point scale) provides more fine-grained descriptions than TOEFL (5-point scale).
Scoring Strategy. In ASAP dataset, the predicted score $\hat{y}$ is rounded to integers. In TOEFL11 dataset, $\hat{y}$ is first scaled to $[1,5]$ (Blanchard et al., 2013), then mapped to low/medium/high with respect to the thresholds of 2.25 and 3.75.

### v2

# Unleashing Large Language Models’ Proficiency in Zero-shot Essay Scoring

@@ Abstract
Advances in automated essay scoring (AES) have traditionally relied on labeled essays, requiring tremendous cost and expertise for their acquisition. Recently, large language models (LLMs) have achieved great success in various tasks, but their potential is less explored in AES. In this paper, we show that our zero-shot prompting framework, Multi Trait Specialization (MTS), elicits LLMs’ ample potential for essay scoring. In particular, we automatically decompose writing proficiency into distinct traits and generate scoring criteria for each trait. Then, an LLM is prompted to extract trait scores from several conversational rounds, each round scoring one of the traits based on the scoring criteria. Finally, we derive the overall score via trait averaging and min-max scaling. Experimental results on two benchmark datasets demonstrate that MTS consistently outperforms straightforward prompting (Vanilla) in average QWK across all LLMs and datasets, with maximum gains of 0.437 on TOEFL11 and 0.355 on ASAP. Additionally, with the help of MTS, the small-sized Llama2-13b-chat substantially outperforms ChatGPT, facilitating an effective deployment in real applications.
Unleashing Large Language Models’ Proficiency
in Zero-shot Essay Scoring
Sanwoo Lee1,2 Yida Cai1,3 Desong Meng1,2 Ziyang Wang1,3 Yunfang Wu1,2thanks:   Corresponding author. 1National Key Laboratory for Multimedia Information Processing, Peking University 2School of Computer Science, Peking University 3School of Software and Microelectronics, Peking University {sanwoo, wuyf}@pku.edu.cn, {caiyida, 2100013162, wzy232303}@stu.pku.edu.cn
@@ 1 Introduction
Automated essay scoring (AES) aims at evaluating and scoring essays with machine learning (Dikli, 2006). AES is a promising alternative to costly and laborious human assessment, greatly resolving rater fatigue and inter-rater inconsistency. AES systems have been widely deployed in classroom settings (Dikli and Bleyle, 2014) and high-stakes tests such as TOEFL (Attali and Burstein, 2006).
@@
Figure 1: Comparison of our MTS zero-shot prompting framework and Vanilla baseline across different types of LLMs and datasets, measured on average QWK.
@@
Previous studies highly matched human ratings via developing supervised models tailored to a specific prompt (Yannakoudakis et al., 2011; Taghipour and Ng, 2016; Dong et al., 2017), assuming essays from the train and test sets belong to the same prompt. However, prompt-specific models struggled when confronted with essays written for unseen prompts (Jin et al., 2018; Cozma et al., 2018). Hence AES is progressing towards reflecting more real-world scenarios, exemplified by cross-prompt approaches (Cao et al., 2020; Jiang et al., 2023b) which reinforce domain transferability of the supervised models.
@@
Recently, advances in large language models (LLMs) (Brown et al., 2020; Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023) have led to a paradigm shift in which LLMs excel across a wide range of downstream tasks via zero-shot or few-shot instructions (Yuan et al., 2023; Zhang et al., 2023a). In many cases, careful prompt design plays a crucial role in unlocking LLMs’ potential. For instance, chain-of-thought (CoT) prompting (Wei et al., 2022; Kojima et al., 2022) improves LLMs’ performance on complex reasoning benchmarks by externalizing the reasoning process.
@@
The development of LLM-based chatbots aligned with human preferences (Ouyang et al., 2022; Rafailov et al., 2023) has given rise to zero-shot AES, allowing us to move beyond the cross-prompt setting. However, leveraging LLMs for zero-shot AES is less explored, in contrast to proliferating studies harnessing LLMs to serve as an evaluation metric for machine-generated text (Chiang and Lee, 2023; Zheng et al., 2023; Liu et al., 2023). Initial works of zero-shot AES (Mizumoto and Eguchi, 2023; Yancey et al., 2023) prompt LLMs to assign the overall score within a single step, which demonstrates suboptimal agreement with human raters.
@@
In this paper, we present MTS (Multi Trait Specialization), a zero-shot prompting framework to elicit essay scoring capabilities in LLMs, inspired by supervised models that explicitly predict trait scores and improve the overall scoring (Ridley et al., 2021; Kumar et al., 2022; Do et al., 2023). In particular, we exploit ChatGPT (OpenAI, 2022) to decompose the writing quality into multiple traits and generate scoring criteria for each trait. Next, an LLM engages in several rounds of conversation, each round evaluating with respect to one of the traits. During the conversation, the LLM is instructed to retrieve quotes to provide faithful evaluation on the essay, then assign a score based on the given scoring criteria. Finally, the overall score is derived by averaging and min-max scaling the trait scores, in combination with the outlier clipping mechanism.
@@
We evaluate MTS on ASAP (Hamner et al., 2012) and TOEFL11 (Blanchard et al., 2013) with different LLMs, including ChatGPT, Llama 2 (Touvron et al., 2023) and Mistral 7b (Jiang et al., 2023a). We take the Vanilla approach as a primary baseline which asks LLMs to produce rationales followed by an overall score. As illustrated in Figure 1, MTS consistently outperforms Vanilla in average Quadratic Weighted Kappa (QWK) across all combinations of LLMs and datasets, with maximum gains of 0.437 ($0.025\rightarrow 0.462$) on TOEFL11 and 0.355 ($0.205\rightarrow 0.560$) on ASAP. In addition, the small-sized Llama2-13b-chat substantially surpasses ChatGPT with the help of MTS, enabling a more effective deployment.
@@
  • Our framework is free from training as well as labeling essays, which can be readily applied to new essay prompts (domains).
  • We automatically decompose the AES task into more specific subtasks with respect to diverse traits, thereby significantly boosting the agreement with human raters.
  • We utilize min-max scaling with outlier clipping, effectively addressing LLMs’ scoring bias and contributing to robust performance.
  • MTS achieves promising results, largely exceeding Vanilla, also outperforming ChatGPT with the small-sized Llama2-13b-chat.
Figure 2: Illustration of the prompt for multi trait decomposition used for ASAP. The contents to be filled are denoted in red. See Appendix A for the templates used for ASAP and TOEFL11.
@@ 2 Method
We formalize the definition of zero-shot AES as follows: Given a dataset consisting of unlabelled essays $\mathcal{D}=\{x^{(i)}\}_{i=1}^{N_{\mathcal{D}}}$, the goal is to output an overall score $\hat{y}^{(i)}$ for every essay $x^{(i)}$ from a set of predefined scores $\mathcal{Y}$ where $\hat{y}^{(i)}\in\mathcal{Y}$.
@@ 2.1 Multi Trait Decomposition
We address this issue by decomposing the writing proficiency into several key traits and defining trait-specific scoring criteria which will be fixed during the assessment (details in Section 2.2), contributing to a consistent scoring behavior of LLM. We automate this process via instructing ChatGPT to condense the rubric guidelines used by human raters into several key traits of writing quality and generate scoring criteria for each trait, using the prompt outlined in Figure 2. This procedure ensures isolating trait-specific scoring criteria from rubric guidelines that mix multiple traits as a whole. An example of the generation result is depicted in the left part of Figure 3 (see Appendix B for more results).
@@ 2.2 Trait Specialization
In Step 1, leveraging trait-specific conversations not only simplifies the AES task as a form of subproblem decomposition but also prevents evaluations on each trait from being influenced by the other traits. In Step 2 and 3, the quote retrieval task is followed by the main scoring task, as depicted in the middle part of Figure 3. The quote retrieval task allows LLM to adhere to the details of the essay and avoid producing generic evaluation, whereas the scoring task transforms the verbal evaluation into the score based on the predefined scoring criteria.
@@ 2.3 Trait Aggregation and Scaling
Parsing the output of Section 2.2 yields trait scores $\{\hat{y}_{j}^{(i)}\}_{j=1}^{N_{T}}$ for an essay $x^{(i)}$, where $N_{T}$ represents the number of predefined traits. The trait scores should be transformed to an overall score $\hat{y}^{(i)}$ that falls under the target score range which may vary across different prompts (as in Table 1).
Based on the trait scores, we devise a simple yet effective trait aggregation and scaling strategy. First, the trait scores are aggregated by taking their average, and the outliers among the averaged scores are clipped using Q1 and Q3, i.e. the first and the third quartiles of the averaged scores, that is, the clipped score $\hat{y}_{agg}^{(i)}$ is computed as follows:
@@
where $\hat{y}_{min}=\min\limits_{i}\hat{y}_{agg}^{(i)}$ and $\hat{y}_{max}=\max\limits_{i}\hat{y}_{agg}^{(i)}$. In this way, the clipping alleviates the sensitivity of min-max scaling to the outliers.
@@ 3.1 LLMs
To verify the effectiveness of our proposed method is universal across LLMs, we choose different types of LLMs which are not variants of one another: ChatGPT, Llama 2 and Mistral 7b. In detail, we use their instruction-tuned models which are gpt-3.5-turbo-0613, Llama2-7b-chat, Llama2-13b-chat and Mistral-7B-Instruct-v0.2. The temperature is set to $0.1$ for all LLMs and the repetition penalty is set to $1.1$ for all LLMs but ChatGPT. Other hyperparameters for sampling follow the defaults. Experiments are run once with a fixed random seed due to limits of computational resources.
@@ 3.2 Datasets and Evaluation Metric
We conduct experiments on two datasets, ASAP [^1] https://www.kaggle.com/c/asap-aes/data and TOEFL11. ASAP (Automated Student Assessment Prize) comprises 12,978 essays provided by students spanning in grade level from 7 to 10. The essays were written in response to 8 prompts of varying genres and score ranges. TOEFL11 consists of 12,100 essays written by test takers of TOEFL iBT which measures the academic English proficiency of non-native English speakers. The statistics of the two datasets are shown in Table 1.
For ASAP, 10% of the essays from each prompt are randomly sampled for testing. There is no significant difference in the average essay score between the sample and the population in all prompts (Z-test, $\alpha=0.05$). For TOEFL11 dataset, we adopt the test split from the 2013 Native Language Identification Shared Task (Tetreault et al., 2013), which consists of 1,100 essays collected from 8 prompts.
We use Quadratic Weighted Kappa (QWK) to measure the agreement between groundtruth scores and predicted scores. QWK is commonly used in AES research (Taghipour and Ng, 2016; Dong et al., 2017; Cao et al., 2020; Xie et al., 2022).
@@
Dataset LLM Method P1 P2 P3 P4 P5 P6 P7 P8 Avg.
ASAP ChatGPT Vanilla 0.032 0.220 0.476 0.597 0.479 0.637 0.289 0.527 0.407
@@
TOEFL11 ChatGPT Mizumoto and Eguchi (2023) 0.308 0.165 0.252 0.182 0.181 0.336 0.318 0.318 0.258
Vanilla 0.215 0.240 0.337 0.332 0.227 0.306 0.237 0.306 0.275
@@
Llama2-7b-chat Mizumoto and Eguchi (2023) 0.009 0.047 0.085 0.140 0.133 0.027 0.032 0.023 0.062
Vanilla $0$0.007 0.026 -0.006 0.041 0.015 0.111 0.020 0.025
@@
Llama2-13b-chat Mizumoto and Eguchi (2023) 0.125 0.132 0.400 0.130 0.462 0.176 0.113 0.123 0.208
Vanilla 0.196 0.156 0.285 0.268 0.165 0.329 0.249 0.257 0.238
@@
Mistral-7b-instruct Mizumoto and Eguchi (2023) 0.227 0.218 0.383 0.350 0.222 0.129 0.132 0.196 0.232
Vanilla 0.486 0.259 0.355 0.344 0.431 0.456 0.286 0.383 0.375
@@
Table 2: Zero-shot evaluation results in QWK. P1-8 denotes Prompt 1-8. The best measures in each LLM are in bold. Negative value indicates that the predictions are worse than random.
@@ 3.3 Implementation Details
Scoring Criteria. As outlined in Section 2.1, ChatGPT generates scoring criteria based on the rubric guidelines for human raters. For ASAP, we use the dataset’s original rubric guidelines for each prompt. For TOEFL11, we follow Mizumoto and Eguchi (2023) and choose IELTS Task2 Writing Band Descriptor as the rubric guidelines instead of TOEFL Independent Writing Rubrics since the former provides more fine-grained descriptions. Note that each prompt in ASAP has a distinct scoring criteria whereas all prompts in TOEFL11 share the same scoring criteria.
Scoring Strategy. For ASAP, the predicted score $\hat{y}$ is rounded to integers. For TOEFL11, $\hat{y}$ is first scaled to $[1,5]$ (Blanchard et al., 2013), then mapped to low/medium/high with respect to the thresholds of 2.25 and 3.75.
Figure 4: Comparison between MTS (Mistral-7b-instruct) and supervised SOTA (NPCR).